-2x+3[x-(-2x+5)]=-(15+2x)
Found 2 solutions by fractalier, fatump:
Answer by fractalier(2101)

-2x + 3[x - (-2x + 5)]= -(15+2x)
-2x + 3(3x - 5) = -15 - 2x
-2x + 9x - 15 = -15 - 2x
7x - 15 = -15 - 2x
9x = 0
x = 0

Answer by fatump(1)